# Embodied Agents Meet Personalization: Exploring Memory Utilization for Personalized Assistance

### v1

@@ D.3 Memory Quality Analysis
D.4 Success and Error Case Analysis
Tasks require object semantics knowledge.
Success cases demonstrate that agents can effectively reference personalized object attributes from episodic memory, correctly identifying and applying the specific information needed for task completion. (A.1) However, we observed distinct error patterns when agents needed to utilize this type of knowledge: missed personalization cues (B.1), where agents failed to recognize the need to access personalized knowledge; hallucinations (B.2), where agents fabricated non-existent attributes; and memory recall failures (B.3), where agents were unable to locate relevant information despite its presence in the provided context.
Tasks require user patterns.
D.5 Details of Discussion Section Experiments
Experiment setup details for Section 6.1.
We sampled 30 episodes by selecting 10 episodes from each of three scenes, with careful balancing across task types and difficulty levels informed by our preliminary experimental results. To minimize the influence of erroneous trajectories, we provided GPT-4o-generated gold memory as the given episodic memory, thereby reducing the impact of noisy interaction histories. We also used GPT-4o to generate summaries of the action trajectories, including user instructions, to facilitate compact memory representations. For each evaluation, one exemplar shot was given to the agent, and top-$k$ memory entries were retrieved based on the current query, where $k=10$ for frontier models (GPT-4o and Qwen-2.5-72B) and $k=7$ for smaller models (LLaMA-3.1-8B and Qwen-2.5-7B), considering context length limitations. The prompts used for these experiments are provided in Appendix E.3.
Experiment setup details for Section 6.2.
@@ D.5 Details of Discussion Section Experiments
We sampled 30 episodes, selecting 10 episodes from each of three scenes, and tailored them by modifying the personalized knowledge statements and reference components. To mitigate potential bias, all episodes were jointly created and validated by two authors. For each model, we first generated memory traces during the memory acquisition stage and then evaluated memory usage during the utilization stage. Table 8 provides representative examples of the ambiguous instruction pairs used. In each case, the user rephrases the original request to refer indirectly to previously established personalized knowledge, simulating natural variability in real-world human-agent communication.
@@
Appendix E Prompts
E.1 Prompts for Dataset Generation
E.2 Prompts for Agent
E.3 Prompts for Discussion
Appendix F Extended Related Work
Recent research has increasingly emphasized the integration of memory mechanisms into large language model (LLM) agents to support long-term reasoning, planning, and personalization. Park et al. [36] propose Generative Agents, which simulate human-like behavior by maintaining a memory stream of past experiences in natural language. This enables agents to reflect, retrieve, and plan based on their individual histories. Similarly, Xu et al. [53] introduce A-Mem, a dynamic memory system inspired by Zettelkasten, which structures memory as evolving and interconnected notes that the agent can generate, retrieve, and update over time, supporting agentic autonomy and adaptability.
Personalization in dialogue agents has also been explored through memory-enhanced architectures. Li et al. [27] present a personalized dialogue agent that leverages both short-term and long-term memories to maintain user-specific context across sessions, significantly improving response consistency and contextual relevance. Wang et al. [46] propose the Self-Controlled Memory (SCM) framework, where an agent dynamically decides when and what to store or retrieve from memory, leading to improved coherence and knowledge retention over extended interactions.
Zhong et al. [60] introduce MemoryBank, a structured long-term memory module that enhances LLMs by storing and retrieving relevant interaction history to support consistent and personalized user responses across multiple turns. Das et al. [12] present Larimar, a framework that integrates episodic memory control into LLMs, enabling selective recall and forgetting to improve memory scalability and privacy-aware reasoning.
More broadly, Packer et al. [35] conceptualize memory management as an OS-level abstraction with MemGPT, enabling an LLM to autonomously manage and interact with its internal and external memory hierarchies. This work highlights how memory can act as a core architectural layer to enable scalable, autonomous agents capable of long-horizon tasks and continuous learning.
Appendix G License
For our implementation and evaluation, we use Huggingface library [^4] https://huggingface.co/ and vLLM library. Both libraries are licensed under Apache License, Version 2.0. We have confirmed that all of the artifacts used in this paper are available for non-commercial scientific use.
G.1 License for the Assets
The existing assets used in this research are properly credited and their licenses respected:
G.2 License for the Codes
The existing codes used in this research are properly credited and their licenses respected:

### v2

@@ E.2.1 Memory Quality Analysis
E.2.2 Robustness to Noisy and Partial Memories
Table 9: Model performance across memory acquisition and utilization stages in Memento.
[…]
GPT-4o Gold Baseline 14.5 2481.2 85.3 80.6
Noisy Random 14.6 2759.0 81.5 -3.8 78.6 -2.0
Shuffle 14.4 2450.9 84.0 -1.3 80.6 0.0
Partial Random 14.5 2526.9 84.4 -0.9 80.1 -0.5
w/o Pick 14.6 2555.7 82.9 -2.4 79.1 -1.5
w/o Place 14.5 2569.8 84.4 -0.9 82.6 +2.0
w/o Nav 15.4 2526.8 81.4 -3.9 78.5 -2.1
Experiment setup.
We conduct additional experiments to evaluate the model’s robustness when episodic memories are corrupted or incomplete. As the target model, We leverage GPT-4o, and we organize the experimental setup into two representative degradation types: Noisy Memory and Partial Memory. For Noisy Memory, we consider two conditions: (i) random step injection, where unrelated steps from other scenes (30%) are inserted into the memory trajectory to simulate retrieval or perception noise, and (ii) temporal shuffle, where the step order within an episode is randomly permuted to disrupt temporal coherence. For Partial Memory, we examine two forms of incompleteness: (i) random step removal, where 20% of steps were randomly removed to test agent robustness under incomplete recall, and (ii) action-type-specific removal, where all steps of a given action type (Place, Place, Navigate) are dropped to analyze their impact on planning.
Random noise and missing navigation cues degrade performance.
As shown in Table 9, the model remains robust under most forms of incomplete memory, with only minor degradation. However, two factors consistently cause substantial performance drops: (i) the injection of random, noisy entries and (ii) the absence of navigation-related information. This indicates that the model is more sensitive to misleading or corrupted inputs than to simple omissions, and that navigation cues play a disproportionately critical role in planning.
E.3 Retrieval Strategy Analysis
Effective retrieval of relevant episodic memories is fundamental to successful memory utilization in personalized assistance tasks. To understand how retrieval quality affects agent performance, we analyze both baseline retrieval effectiveness and the impact of advanced retrieval strategies on downstream task outcomes. We first establish baseline missing rates using similarity-based retrieval to determine optimal retrieval parameters, then examine whether advanced retrieval strategies consistently improve both memory retrieval quality and agent task performance. This analysis provides insights into the relationship between retrieval precision and practical task execution in personalized embodied agents.
Table 10: We use an embedding-based retrieval method with the all-mpnet-base-v2 Sentence Transformer (Reimers & Gurevych, 2019).
Retriever Strategy Top-K Missing Rates (%)
Baseline 3 13.9
5 3.5
E.3.1 Baseline Retrieval Analysis.
To assess whether the corresponding episodic memory is successfully retrieved during memory utilization, we first measure the missing rate under a baseline similarity-based retriever.  [^1] We adopt an embedding-based method using the all-mpnet-base-v2 Sentence Transformer (Reimers & Gurevych, 2019). The results show missing rates of 13.9% for top-3 and 3.5% for top-5, indicating that expanding to top-5 nearly eliminates missing cases. Based on this observation, we set top-5 as the default retrieval setup, prioritizing inclusion of the required memory over the risk of distractor entries.
E.3.2 Advanced Retrieval Strategy Analysis
Table 11: Performance comparison across different retrieval strategies.
Model Strategy Top-K Missing Rates (%) Percent Complete $\uparrow$ Success Rate $\uparrow$
GPT-4o Baseline 3 13.9 82.9 79.5
5 3.5 84.7 82.1
Reranker 3 2.5 87.2 84.1
5 1.5 84.3 80.6
Query Expansion 3 9.0 84.9 81.1
5 1.5 86.2 81.1
Qwen-2.5-72b Baseline 3 13.9 75.3 70.1
5 3.5 75.4 70.0
Reranker 3 7.6 73.0 67.2
5 2.1 70.9 64.1
Query Expansion 3 8.5 75.6 69.2
5 4.0 72.6 66.2
Experiment setup.
Building on this baseline, we further examine the effect of advanced retrieval strategies on the instruction following capability. We compared three approaches: (i) the baseline similarity-based retreival, (ii) reranker-based retrieval, which applies an LLM-based pointwise reranker (Sachan et al., 2022; Sun et al., 2023) to an initial pool of 10 candidates before selecting the final top-3 or top-5, and (iii) query expansion-based retrieval, where queries are reformulated before retrieval (Jagerman et al., 2023; Wang et al., 2023b). These strategies are evaluated with both GPT-4o and Qwen-2.5-72b, using the same models to compute reranker scores and produce expanded queries.
Improved retrieval does not ensure improved outcomes.
As shown in Table 11, retrieval strategies affect not only memory inclusion but also downstream task performance. First, applying reranker or query expansion substantially reduce the missing rate, particularly under the top-3 setting, where improvements range from 4.9pp to 11.4pp. This confirms their effectiveness in narrowing the gap between top-3 and top-5 retrieval. Second, performance gains arere not consistent across models. For example, with Qwen, query expansion under the top-5 setting even degraded retrieval, suggesting that the effectiveness of query expansion critically depends on the quality of the model generating the reformulated queries. Finally, except for GPT-4o at top-3, introducing advanced retrieval strategies further reduces downstream performance, likely due to the inclusion of semantically similar but task-irrelevant memories. These results suggest that simply increasing retrieval relevance does not necessarily yield better task performance, particularly in real-world contexts with dense and noisy memory pools.
E.4 Real-World Scenario Analysis
While our main evaluation demonstrates fundamental limitations in memory utilization for personalized assistance, real-world deployment introduces additional complexities that extend beyond our controlled experimental settings. To explore these practical challenges, we conduct three sets of experiments that examine agent behavior under more realistic conditions. First, we evaluate scalability by testing agents on 3-joint memory tasks that require integrating knowledge from multiple episodes. Second, we assess how agents handle ambiguous instructions that indirectly reference personalized knowledge through contextual cues rather than explicit mentions. Finally, we investigate agents’ ability to generalize learned user preferences across different but related contexts, such as applying spatial arrangements from one room to another. These experiments provide insights into additional hurdles for successful real-world deployment of personalized embodied agents.
E.4.1 Real-World Multi-Episode Evaluation
Table 12: Performance results on multiple joint-memory tasks.
Model Task PC (%) SR (%)
GPT-4o joint memory task 86.7 63.9
GPT-4o 3-joint memory task 61.2 22.0
Qwen-2.5-72b joint memory task 68.9 36.1
Qwen-2.5-72b 3-joint memory task 42.4 7.3
In practice, user requirements are rarely limited to information from just one or two episodes, but often require knowledge distributed across multiple episodes. To explore this realistic scenario, we extend our joint-memory task to cases involving three memories, which require agents to effectively integrate three independent memory sources to solve the task. Following the same methodology as our joint-memory task construction, we sequentially concatenated three personalized instructions to generate a total of 41 tasks. We observed substantial performance drops compared to the two-memory joint-memory tasks for both GPT-4o and Qwen-2.5-72b-instruct. This suggests that multi-memory reasoning introduces significant difficulty due to temporal gaps and abstraction challenges. We believe this degradation reflects the true complexity of real-world personalized tasks, and addressing this challenge remains critical for future research.
E.4.2 Agent Behavior under Ambiguous Instructions from Users
Model PC (%) SR (%)
GPT-4o (Baseline) 92.0 90.0
GPT-4o 80.4 73.3
Qwen-2.5-72b (Baseline) 75.1 66.7
Qwen-2.5-72b 59.6 53.3
Experiment setup.
@@ E.4.2 Agent Behavior under Ambiguous Instructions from Users
While Memento focuses on evaluating personalized knowledge grounding with explicit references to prior interactions, real-world human-agent communication often involves ambiguous or indirect references. To explore this challenge, we conducted a proof-of-concept experiment to assess whether current models can interpret ambiguous instructions that indirectly refer to previously encountered personalized knowledge. We created a tailored set of 30 tasks by sampling 10 episodes from each of three scenes and modifying them to include indirect references to personalized knowledge through contextual cues, synonyms, or causal references (e.g., Can you set my afternoon tea time routine? $\rightarrow$ I’m about to enjoy my afternoon tea. Could you set things up as I like them?). To ensure quality and mitigate potential bias, all episodes were jointly created and validated by two authors. For each model, we first generated memory traces during the memory acquisition stage and then evaluated memory usage during the utilization stage. Table 14 provides representative examples of the ambiguous instruction pairs used. In each case, the user rephrases the original request to refer indirectly to previously established personalized knowledge, simulating natural variability in real-world human-agent communication.
Result.
The results, shown in Table 13, reveal the degradation in performance, indicating that handling ambiguous queries remains a key challenge for future personalized embodied agents. We view this as a promising direction for future work, where we plan to systematically extend Memento to evaluate LLM-powered embodied agents’ capabilities under ambiguous and implicit reference scenarios, aiming to better reflect the complexity of real-world human-agent interaction.
@@
E.4.3 Personalized Knowledge Generalization
Experiment setup.
A critical aspect of personalized embodied agent is their ability to generalize learned user preferences across different but related contexts. This includes applying knowledge about a user’s breakfast routine to dinner preparation tasks, or transferring spatial preferences from one room to another. To investigate this capability, we designed generalization tasks that test whether embodied agents can abstract and reuse personalized knowledge as transferable templates. We extended our user pattern tasks by sampling existing scenarios and systematically altering target locations, thereby evaluating whether agents can adapt user preferences to novel but structurally similar situations. To encourage explicit memory utilization, we incorporated contextual cues in the instructions, such as "on [another receptacle] this time," clearly signaling that established preferences should apply in the modified context. Our generalization tasks were constructed using relatively simple base cases that GPT-4o consistently solves in standard settings.
Table 15: Agent performance on personalized knowledge generalization tasks.
Model Task PC (%) SR (%)
GPT-4o Baseline 100.0 100.0
GPT-4o Generalization 92.5 83.3
Qwen-2.5-72b Baseline 75.0 75.0
Qwen-2.5-72b Generalization 85.8 75.0
Result.
As in Table 15, we observed that GPT-4o occasionally fails by defaulting to commonsense reasoning rather than retrieving and adapting the relevant personalized memory. Interestingly, Qwen-2.5-72b demonstrated improved completion rates on these tasks. We attribute this improvement to the explicit instruction cues, which likely reduced ambiguity and guided the model toward memory-based reasoning. While these preliminary experiments provide initial insights into generalization capabilities, they represent an important step toward understanding how agents can flexibly apply personalized knowledge. Future work could build on these initial observations to better understand how agents transfer personalized knowledge across different contexts.
Appendix F Prompts
F.1 Prompts for Dataset Generation
F.2 Prompts for Agent
F.3 Prompts for Memory Design Experiment
Table 16: Episode metadata for the example task.
Field Value
episode_type user_pattern
subtype preference
task Could you decorate the bedroom as I like?
task_percent_complete 1.0
task_state_success 1.0
Table 17: Trajectory for the example task.
# Thought Action Result / State $\Delta$
1 Since there are no objects found yet, I should start by exploring the bedroom to locate any objects that can be used for decoration. Explore
(bedroom_1)
No objects found
2 Since no objects were found in bedroom_1, I should explore another room to locate objects that can be used for decoration. I will explore the living room next, as it is likely to contain decorative items. Navigate
(to living_room_1)
Location changed
3 I have found a plant container and a statue on the shelves in the living room, which can be used for decoration in the bedroom. I will start by picking up the plant container from the shelves. Pick
(plant_container_1)
plant_container_1 held
4 Now that I am close to the shelves, I will attempt to pick up the plant container to move it to the bedroom for decoration. (continue) Success
5 I have successfully picked up the plant container. Now, I need to navigate to the bedroom to place it for decoration. Navigate
(to bedroom_1)
Arrived in bedroom_1
6 I have successfully navigated to bedroom_1. Now, I need to find a suitable location to place the plant container for decoration. I will place it on the chest of drawers in the bedroom. Place
(plant_container_1
on chest_of_drawers_54)
Object placed
7 Now that I am close to chest_of_drawers_54, I will place the plant container on it for decoration. (continue) Confirmed placement
8 I have successfully placed the plant container on chest_of_drawers_54 in bedroom_1. Now, I need to return to the living room to pick up the statue from the shelves for further decoration in the bedroom. Navigate
(to living_room_1)
Location changed
9 I am back in the living room. I need to navigate closer to the shelves to pick up the statue for further decoration in the bedroom. Positioning for pick
10 Now that I am close to the shelves, I will attempt to pick up the statue to move it to the bedroom for decoration. Pick
(statue_0)
statue_0 held
11 I have successfully picked up the statue. Now, I need to navigate to the bedroom to place it for decoration. Navigate
(to bedroom_1)
Location changed
12 I have successfully navigated to bedroom_1. Now, I need to find a suitable location to place the statue for decoration. I will place it on the chest of drawers next to the plant container. Positioning
13 Now that I am close to chest_of_drawers_54, I will place the statue next to the plant container for decoration. Place
(statue_0 on
chest_of_drawers_54
next_to plant_container_1)
Object placed
Table 18: Evaluation functions for the example task.
# Function Args Description
1 is_on_top object_handles=[Cole_Hardware_Electric_Pot_Assortment_55_:0000], receptacle_handles=[bedroom_chest_of_drawers_handles],
number=1
Plant container is on drawer
2 is_on_top object_handles=[Dog_Statue_4_:0000],
receptacle_handles=[bedroom_chest_of_drawers_handles],
number=1
Statue is on drawer
3 is_next_to entity_handles_a=[Cole_Hardware_Electric_Pot_Assortment_55_:0000], entity_handles_b=[Dog_Statue_4_:0000],
number=1,
l2_threshold=1.0
Plant and statue are next to each other
Table 19: Episode metadata for the example task.
Field Value
episode_type user_pattern
subtype routine
task Could you set up the candle for a cozy evening?
task_percent_complete 0
task_state_success 0
Table 20: Trajectory for the example task.
# Thought Action Result / State $\Delta$
1 Since there are no objects found yet, I should start by exploring the tv room to locate a candle that can be set up for a cozy evening. Navigate
(to tv_1)
Location changed
2 I have found a candle, candle_0, on table_34 in the tv room. To set it up for a cozy evening, I will first need to pick it up from the table. Pick
(candle_0)
candle_0 held
3 Now that I am close to table_34, I will attempt to pick up the candle, candle_0, to set it up for a cozy evening. (continue) Success
4 I have successfully picked up the candle, candle_0. To set it up for a cozy evening, I will place it on a suitable surface in the tv room, such as the tv table, table_14. Place
(candle_0 on table_14)
Object placed
5 I am now close to table_14 in the tv room. I will place the candle, candle_0, on this table to set it up for a cozy evening. (continue) Failed to meet task constraints
Table 21: Evaluation functions for the example task.
# Function Args Description
1 is_on_top object_handles=[candle], receptacle_handles=[chair_handles],
number=1
Candle must be on chair
2 is_on_top object_handles=[candle], receptacle_handles=[table_handles],
number=1
Candle must be on table (after chair)
3 Constraint TemporalConstraint: Placing the candle on the chair before placing it on the table.
Figure 12: Example of a user profile memory graph representing object semantics knowledge.
Figure 13: Example of a user profile memory graph representing user pattern knowledge.